# Attention Is All You Need for LLM-based Code Vulnerability Localization

### v1

@@ 3.2 Exploratory Experiments and Key Idea
Specifically, we have formulated four hypotheses to enhance vulnerability detection in code. First, we hypothesize that accurate localization of a vulnerable line of code necessitates the model’s focused attention on that specific line. Given the challenges posed by extended contexts, vulnerable lines may be overlooked. To address this, our second hypothesis proposes manually intensifying the model’s focus on identified vulnerable lines to improve detection accuracy. Thirdly, we posit that a genuinely vulnerable line will inherently exert a significant influence on the model’s output. However, a major challenge is the inability to pre-identify which lines warrant this increased attention due to unknown vulnerability statuses. Intuitively, we suggest directing the model to scrutinize every line of code, irrespective of its vulnerability status. This approach hinges on our fourth hypothesis: intensifying the model’s focus across all code lines is unlikely to lead to incorrect evaluations, and focusing on non-vulnerable code will not adversely affect the attention dynamics.
@@
[backgroundcolor=blue!2]
Observation (I).The lines identified as vulnerable by the model are consistently highlighted in the attention distribution. However, insufficient attention may prevent the model from effectively detecting the vulnerability. As a result, the output may be prone to a high false positives.
EE-II: We then explored how to direct the model’s attention toward the lines containing vulnerabilities. One straightforward approach is to modify the prompt to guide the model’s focus to a specific line. By instructing the model to “pay attention to line 70”, we successfully shifted its focus. The attention results, shown in Figure 3 (b) and (d), indicate a significant increase in attention on line 70.
@@
[backgroundcolor=blue!2] Observation (II). By adjusting the prompt, we can direct the model to focus on a specific line, effectively altering its attention.
@@
[backgroundcolor=blue!2] Observation (III). By increasing attention on a vulnerable line, we can make it easier for the model to identify it as vulnerable.
@@
[backgroundcolor=blue!2] Observation (IV). When a vulnerable line is highlighted, certain attention patterns show significant increases, which do not appear for non-vulnerable lines.
Key Idea: As outlined in Observations (I)-(IV), LLMs tasked with vulnerability localization in extensive contexts frequently miss the actual vulnerable lines (O-I). By strategically modifying the model’s prompts, we can enhance the model’s focus on these vulnerable lines (O-II), thereby refining its capability to accurately localize vulnerabilities (O-III). Furthermore, directing the model’s attention towards non-vulnerable lines does not adversely affect its overall attention allocation (O-IV). Therefore, our idea is that we can systematically iterate over each line of code, individually highlighting them to ensure comprehensive coverage of all potential vulnerabilities. This method is effective because highlighting a vulnerable line results in a marked increase in the model’s attention, whereas non-vulnerable lines exhibit minimal changes. By monitoring these attention shifts, we can precisely identify and confirm the presence of vulnerabilities, thus significantly improving the accuracy of vulnerability localization.
3.3 Threat Model
Scope and Assumptions: We now discuss the scope and assumptions of our work. The primary goal is to localize the vulnerabilities within a program, guided by the following requirements. First, there are indeed numerous vulnerability localization tools in the software security domain, but our tool specifically focuses on those powered by LLMs, recognizing the immense potential they offer (e.g., ). Second, this paper focuses on the task of vulnerability localization within a single file. We do not address vulnerabilities that span across multiple files, such as those involving inter-file code dependencies or cross-file data flows. For example, vulnerabilities related to privacy leakage often occur when sensitive data is collected in one file, processed in another, and then improperly exposed or transmitted in yet another file. This is out of our focus.
@@ 3.3 Threat Model
4 Design of LOVA
Figure 4: Overview of LOVA. For vulnerable code, the process begins by highlighting each line individually to generate a base prompt and a set of highlighted prompts. These prompts are then processed through the LLM’s prefill phase to obtain their respective attention maps. For each highlighted attention map, the difference with the base attention map is computed, and the result is flattened to form an attention vector corresponding to each highlighted line. These attention vectors are then classified by a Bi-LSTM classifier to determine whether each line contains a vulnerability, ultimately yielding a suspicious score for each line.
In this section, we now discuss the design details of LOVA. As shown in Figure 4, LOVA consists of three primary stages, i.e., code line highlight, attention calculation, and vulnerability line localization. Specifically, in each of these steps, we first outline the challenges (C-1, C-2, C-3) encountered, followed by the corresponding solutions (S-1, S-2, S-3).
@@ 4 Design of LOVA
4.1 Code Line Highlight
In this step, LOVA generates two versions of the input based on the vulnerable program, both of which will be processed by the LLM in subsequent steps. The first version is a direct copy of the vulnerable program with the template prompt inserted, but none of the lines are highlighted (Step ❶). The second version highlights every line of the vulnerable program (Step ❷). To create the second version, LOVA systematically examines each line of code, sequentially applying highlights. This highlighting is intended to direct the LLM’s focus to specific lines. By highlighting a code line, the attention distribution across the prompt is altered (Step ❸).
(C-I) - Inappropriate prompt may mislead LLM: At first glance, highlighting a specific line may seem straightforward, with several methods available to achieve this. For instance, comments like //Pay attention to this or markers such as /*FOCUS_BEGIN*/ and /*FOCUS_END*/ can be inserted into the code to draw attention to a specific line [22, 21]. Another approach is to mask unimportant sections of the input, allowing the model to focus exclusively on the critical parts [20].
@@ 4.1 Code Line Highlight
[backgroundcolor=green!2] (S-I) - Highlight lines using their index. Our approach is straightforward yet effective. Instead of adding extra content that might confuse the LLM, we directly highlight specific lines by specifying their line index. This method accurately enhances attention on the highlighted lines and improves the LLM’s success rate in locating them. Additionally, the differences between lines are significant and easily analyzable, providing clear, quantifiable reference points for future analysis.
Next, the key question is how to construct an effective prompt based on (S-I). As depicted in Figure 1, we have implemented the following steps: First, we prepend line indices to the code and instruct the model to ”pay attention to” a specific index, thereby enhancing focus on that particular line. These indices are based on absolute positions, rather than relative ones, as LLMs have difficulty with relative positioning [19]. Second, our objective is to direct the model’s attention to the highlighted line and prompt it to determine whether that line contains a vulnerability, without introducing any prior assumptions. To this end, we instruct the LLM with the primary objective of identifying potential vulnerabilities in the entire code, while focusing specifically on the correctness of the highlighted line, rather than only judging the correctness of that line. Finally, to facilitate future analysis, we minimize the inclusion of non-code lines in the prompt. Since highlighting primarily impacts the attention directed toward the highlighted line and the instructions, reducing non-code lines helps streamline subsequent analyses, making it easier to track the model’s attention shifts and performance.
@@
((a)) Highlighting the 8th line of code
Figure 5: 1 out of 10 highlighted prompt examples for one code
4.2 Attention Calculation
@@
(C-2) Impracticality of analyzing large attention tensors: In the original attention output, a tensor with the shape (num_tokens, num_tokens, num_layers, num_heads) is generated, representing the attention scores calculated for every pair of tokens across all layers and heads of the model. This structure captures the intricate relationships between tokens and provides valuable insights into how attention is distributed throughout the network:
@@
  • The second num_tokens represents the relationship of each token in the code (or text) with all other tokens that follow it.
@@
  • num_heads refers to how many independent attention heads exist in each layer. Each head focuses on different aspects of the code to help the model better understand its structure and semantics.
@@
  • Multi-Layer Attention Preservation (Step ➁). The attention from all layers is retained because each decoder layer possesses a unique capacity to capture different semantic and relational features. For instance, lower layers may concentrate on fundamental patterns such as syntax, while higher layers tend to capture more abstract concepts, like the overall meaning of the code [14]. By preserving attention across all layers, we ensure the model captures both basic and complex information, enriching the representation.
@@
  • Line-Level Attention Summation (Step ➃). The attention values of all tokens in each line of the prompt are added together to represent the overall attention for that line. This process results in a matrix of size (num_lines, num_layers). We refer to this matrix as the LayerwiseAttnMat. The key idea here is that attention is not evenly spread across all tokens—some tokens get more focus than others. By summing the attention for each line, we simplify the data while still capturing the most important information, since many tokens may not show significant differences in attention.
@@
4.3 Vulnerability Line Localization
@@
(C-3) Precise Cross-Language Vulnerability Localization. We found that attention patterns for code or text within a single programming language (e.g., Java, C) often exhibit similarities. These patterns tend to reflect semantic information related to vulnerabilities at specific lines and layers in the code. Therefore, summing the attentions at these positions on the VulnAttnMat can yield a corresponding suspicion score (which is a numerical value assigned to data point that quantifies how “suspicious” or anomalous it is) that highlights which sections of the code are more likely to contain vulnerabilities. However, vulnerabilities differ greatly across languages. For instance, in network protocols written in C, a buffer overflow might occur during packet processing, whereas in a web application, the vulnerability might involve cross-site scripting in JavaScript. These varying contexts mean that attention patterns that indicate a vulnerability in one language may not be applicable to other languages. Additionally, relying solely on these summed attention values provides a suspicion score, which only ranks how suspicious a line of code is compared to others. While this can prioritize certain areas for review, it does not provide a probability for each line being vulnerable, meaning it lacks precision. For effective vulnerability localization, it’s crucial to assess the actual probability of each line containing a vulnerability, rather than just ranking the lines by suspicion.
@@
[backgroundcolor=green!2] (S-III) Language-aware deep learning-based vulnerability localization. We chose to employ a deep learning approach to classify the attention matrix and determine whether the lines contain vulnerabilities. This method can leverage complex patterns and relationships within the data to provide a more accurate and probabilistic assessment. Unlike simply ranking lines by suspicion score, a deep learning model can learn to identify specific vulnerability signatures across different languages, improving precision in vulnerability localization.
@@
To address this, LOVA implements a learning approach that integrates sequential information by employing a Bi-LSTM model, which effectively captures these contextual dependencies: First, The identification of anomalies of VulnAttnMat is context-dependent. For example, when inputting three lines of code, we generate three corresponding VulnAttnMat: $A$, $B$ and $C$. If $A$ significantly deviates from $B$ and $C$, it suggests that $A$ may be the most suspicious. To effectively model this, a sequence-to-sequence approach is necessary, with LSTM [34] and Transformer being the most common models for such tasks. Second, given that the features of anomalies in our case are relatively distinct, a straightforward LSTM model is sufficient. However, we opted for a Bi-LSTM [33] to leverage its ability to capture bidirectional dependencies, as standard LSTM only processes information in one direction. This is important because the detection of anomalies is influenced by both the preceding and succeeding contexts, rather than solely relying on prior context.
@@
where $N$ is the number of lines, $y_{i}$ is the true label of the $i$-th line, and $\hat{y}_{i}$ is the predicted probability of the line containing a vulnerability. During the evaluation phase, the flattened vectors of the VulnAttnMat are input into the model to compute suspicious scores for each line. These scores are subsequently ranked in descending order to inform the vulnerability localization process, with scores greater than 0.5 being considered as vulnerable lines.
5 Evaluation
@@
    How does the design of LOVA impact the results (Abalation Study)?
5.1 Experiment Setup
Dataset. Our dataset, comprises vulnerability benchmarks from C/C++ projects (Big-Vul), Solidity smart contracts (SmartFix), and multi-language CVE records (CVEFixes). This allows for a comprehensive evaluation of LOVA’s performance across diverse contexts. Specifically:
  • Big-Vul [9]: A C/C++ code vulnerability benchmark collected from 384 open-source GitHub projects. Big-Vul includes both vulnerable code snippets and their corresponding fixing commits. Vulnerabilities are annotated at function context level, categorizing the types of vulnerabilities and associated commits. Big-Vul contains a total of 11,823 vulnerable functions.
  • SmartFix [8]: A Solidity code vulnerability benchmark collected from multiple sources. SmartFix covers five common types of smart contract vulnerabilities: Integer Overflow, Reentrancy, Ether Leak, Suicidal, and incorrect use of ‘tx.origin‘. Each vulnerability is annotated with its corresponding line numbers. SmartFix contains a total of 361 contract files.
  • CVEFixes [7]: A multi-language code vulnerability benchmark that is automatically collected and curated from CVE records in the public U.S. National Vulnerability Database (NVD). CVEFixes collects both vulnerabilities and the commits that fix them, including 5,365 CVEs and 50,322 methods. CVEFixes encompasses multiple languages, such as C, C++, PHP, Python, Java, and JavaScript. From these, we selected the three most commonly used languages, C, Python and Java, as benchmarks.
Big-Vul provides vulnerability data at the function context level, CVEFixes provides vulnerability data at both the function and file context levels, and SmartFix provides vulnerability data at the contract context level. To evaluate LOVA’s performance across different contexts, we conduct evaluations at the function, file, and contract context levels. Additionally, we filter out contexts that exceed 4,000 tokens to comply with the input constraints of LLMs. In total, seven datasets were created, as shown in Table I. Among them, the C language datasets Big-Vul and CVEFixes-C each contain 1,000 instances. For Java, there are two datasets: CVEFixes-J-M and CVEFixes-J, corresponding to method-level and file-level contexts, respectively. In Python, the CVEFixes-P-M and CVEFixes-P datasets represent method-level and file-level contexts, respectively. Additionally, Solidity has one dataset, SmartFix, which focuses on contract-level context.

### v2

@@ 3.2. Exploratory Experiments and Key Idea
Specifically, we have formulated four hypotheses to enhance vulnerability localization in code. First, we hypothesize that accurate localization of a vulnerable line of code necessitates the model’s focused attention on that specific line. Given the challenges posed by extended contexts, vulnerable lines may be overlooked. To address this, our second hypothesis proposes manually intensifying the model’s focus on identified vulnerable lines to improve detection accuracy. Thirdly, we posit that a genuinely vulnerable line will inherently exert a significant influence on the model’s output. However, a major challenge is the inability to pre-identify which lines warrant this increased attention due to unknown vulnerability statuses. Intuitively, we suggest directing the model to scrutinize every line of code, irrespective of its vulnerability status. This approach hinges on our fourth hypothesis: intensifying the model’s focus across all code lines is unlikely to lead to incorrect evaluations, and focusing on non-vulnerable code will not adversely affect the attention dynamics.
@@
[backgroundcolor=orange!4]
Observation (I). The lines identified as vulnerable by the model are consistently highlighted in the attention distribution. However, insufficient attention may prevent the model from effectively detecting the vulnerability. As a result, the output may be prone to high false positives.
EE-II: We then explored how to direct the model’s attention toward the lines containing vulnerabilities. One straightforward approach is to modify the prompt to guide the model’s focus to a specific line(i.e., to highlight this line). By instructing the model to “pay attention to line 70”, we successfully shifted its focus. The attention results, shown in Figure 3 (b) and (d), indicate a significant increase in attention on line 70.
@@
[backgroundcolor=orange!4] Observation (II). By adjusting the prompt, we can direct the model to focus on a specific line, effectively altering its attention.
@@
[backgroundcolor=orange!4] Observation (III). By increasing attention to a vulnerable line, we can make it easier for the model to identify it as vulnerable.
@@
[backgroundcolor=orange!4] Observation (IV). When a vulnerable line is highlighted, certain attention patterns show significant increases, which do not appear for non-vulnerable lines.
Key Idea: As outlined in Observations (I)-(IV), LLMs tasked with vulnerability localization in extensive contexts frequently miss the actual vulnerable lines (O-I). By strategically modifying the model’s prompts, we can enhance the model’s focus on these vulnerable lines (O-II), thereby refining its capability to accurately localize vulnerabilities (O-III). Furthermore, directing the model’s attention towards non-vulnerable lines has minimal impact on its overall attention allocation (O-IV). Therefore, our idea is that we can systematically iterate over each line of code, individually highlighting them to ensure comprehensive coverage of all potential vulnerabilities. This method is effective because highlighting a vulnerable line results in a marked increase in the model’s attention, whereas non-vulnerable lines exhibit minimal changes. By monitoring these attention shifts, we can precisely identify and confirm the presence of vulnerabilities, thus significantly improving the accuracy of vulnerability localization.
3.3. Threat Model
Scope and Assumptions: We now discuss the scope and assumptions of our work. The primary goal is to localize the vulnerabilities within a program, guided by the following requirements. First, there are indeed numerous vulnerability localization tools in the software security domain, but our tool specifically focuses on those powered by LLMs, recognizing the immense potential they offer (e.g., CoT (Wei et al., 2022)). Second, this paper focuses on the task of vulnerability localization within a single file. We do not address vulnerabilities that span across multiple files, such as those involving inter-file code dependencies or cross-file data flows. For example, vulnerabilities related to privacy leakage often occur when sensitive data is collected in one file, processed in another, and then improperly exposed or transmitted in yet another file. This is out of our focus. It is important to emphasize that our work focuses on vulnerability localization (Xu et al., 2024; Shen et al., 2021; Yin et al., 2024) rather than vulnerability detection, meaning that all input programs are expected to contain at least one vulnerability. Our approach helps developers quickly pinpoint the location of vulnerabilities when a vulnerability trigger or crash is identified.
@@ 3.3. Threat Model
4. Design of LOVA
Figure 4. Overview of LOVA. For vulnerable code, the process begins by highlighting each line individually to generate a base prompt and a set of highlighted prompts. These prompts are then processed through the LLM’s prefill phase to obtain their respective attention maps. For each highlighted attention map, the difference with the base attention map is computed, and the result is flattened to form an attention vector corresponding to each highlighted line. These attention vectors are then classified by a Bi-LSTM classifier to determine whether each line contains a vulnerability, ultimately yielding a suspicious score for each line.
In this section, we now discuss the design details of LOVA. As shown in Figure 4, LOVA consists of three primary stages, i.e., code line highlight, attention calculation, and vulnerability line localization. Specifically, in each of these steps, we first outline the challenges (C-I, C-II, C-III) encountered, followed by the corresponding solutions (S-I, S-II, S-III).
@@ 4. Design of LOVA
4.1. Code Line Highlight
In this step, LOVA generates two versions of the input based on the vulnerable program, both of which will be processed by the LLM in subsequent steps. The first version is a direct copy of the vulnerable program with the template prompt inserted, but none of the lines are highlighted (Step ❶). The second version highlights each line of the vulnerable program (Step ❷). To create the second version, LOVA systematically examines each line of code, sequentially applying highlights. This highlighting is intended to direct the LLM’s focus to specific lines. By highlighting a code line, the attention distribution across the prompt is altered (Step ❸).
(C-I) Inappropriate prompt may mislead LLM: At first glance, highlighting a specific line may seem straightforward, with several methods available to achieve this. For instance, comments like //Pay attention to this or markers such as /*FOCUS_BEGIN*/ and /*FOCUS_END*/ can be inserted into the code to draw attention to a specific line (Jiang et al., 2023a; Prenner et al., 2022). Another approach is to mask unimportant sections of the input, allowing the model to focus exclusively on the critical parts (Berabi et al., 2024).
@@ 4.1. Code Line Highlight
[backgroundcolor=green!2] (S-I) Highlight lines using their index. Our approach is straightforward yet effective. Instead of adding extra content that might confuse the LLM, we directly highlight specific lines by specifying their line index. This method accurately enhances attention on the highlighted lines and improves the LLM’s success rate in locating them. Additionally, the differences between lines are significant and easily analyzable, providing clear, quantifiable reference points for future analysis.
Next, the key question is how to construct an effective prompt based on (S-I). As depicted in Figure 5, we have implemented the following steps: First, we prepend line indices to the code and instruct the model to ”pay attention to” a specific index, thereby enhancing focus on that particular line. These indices are based on absolute positions, rather than relative ones, as LLMs have difficulty with relative positioning (Zhang et al., 2024a). Second, our objective is to direct the model’s attention to the highlighted line and prompt it to determine whether that line contains a vulnerability, without introducing any prior assumptions. To this end, we instruct the LLM with the primary objective of identifying potential vulnerabilities in the entire code, while focusing specifically on the correctness of the highlighted line, rather than only judging the correctness of that line. Finally, to facilitate future analysis, we minimize the inclusion of non-code lines in the prompt. Since highlighting primarily impacts the attention directed toward the highlighted line and the instructions, reducing non-code lines helps streamline subsequent analyses, making it easier to track the model’s attention shifts and performance.
@@
Figure 5. 1 out of 10 highlighted prompt examples for one code (highlighting line 8)
4.2. Attention Calculation
@@
(C-II) Impracticality of analyzing large attention tensors: In the original attention output, a tensor with the shape (num_tokens, num_tokens, num_layers, num_heads) is generated, representing the attention scores calculated for every pair of tokens across all layers and heads of the model. This structure captures the intricate relationships between tokens and provides valuable insights into how attention is distributed throughout the network:
@@
  • The second num_tokens represents the relationship of each token in the code (or text) with all other tokens that follow it (the first num_tokens).
@@
  • num_heads refers to how many independent attention heads exist per layer. Each head focuses on different aspects of the code’s structure and semantics.
@@
  • Multi-Layer Attention Preservation (Step ➁). The attention from all layers is retained because each decoder layer possesses a unique capacity to capture different semantic and relational features. For instance, lower layers may concentrate on fundamental patterns such as syntax, while higher layers tend to capture more abstract concepts, like the overall meaning of the code (Ma et al., 2024a). By preserving attention across all layers, we ensure the model captures both basic and complex information, enriching the representation.
@@
  • Line-Level Attention Summation (Step ➃). The attention values of all tokens in each line of the prompt are added together to represent the overall attention for that line. This process results in a matrix of size (num_lines, num_layers). We refer to this matrix as the LayerwiseAttnMat. The key idea here is that attention is inherently sparse—not all tokens receive equal focus (Liu et al., 2022, 2023). Some tokens get significantly more attention than others. By summing the attention for each line, we simplify the data while still capturing the most important information, as many tokens contribute minimally to the overall attention distribution.
@@
4.3. Vulnerability Line Localization
@@
(C-III) Precise Multi-Language Vulnerability Localization: We found that attention patterns for code or text within a single programming language (e.g., Java, C) often exhibit similarities. These patterns tend to reflect semantic information related to vulnerabilities at specific lines and layers in the code. Therefore, summing the attentions at these positions on the VulnAttnMat can yield a corresponding suspicion score (which is a numerical value assigned to a data point that quantifies how “suspicious” or anomalous it is) that highlights which sections of the code are more likely to contain vulnerabilities. However, vulnerabilities differ greatly across languages. For instance, in network protocols written in C, a buffer overflow might occur during packet processing, whereas in a web application, the vulnerability might involve cross-site scripting in JavaScript. These varying contexts mean that attention patterns that indicate a vulnerability in one language may not be applicable to other languages. Additionally, these summed attention values serve as a coarse heuristic, providing a suspicion score that only ranks the relative suspiciousness of code lines, but falls short of achieving accurate vulnerability localization. While this can prioritize certain areas for review, it does not provide a probability for each line being vulnerable. For effective vulnerability localization, it’s crucial to assess the actual probability of each line containing a vulnerability, rather than just ranking the lines by suspicion.
@@
[backgroundcolor=green!2] (S-III) Language-aware deep learning-based vulnerability localization. We chose to employ a deep learning approach to classify the attention matrix and determine whether the lines contain vulnerabilities. This method can leverage complex patterns and relationships within the data to provide a more accurate and probabilistic assessment. Unlike simply ranking lines based on a suspicion score obtained by crudely summing attention values, a deep learning model can learn to identify specific vulnerability signatures in different languages, improving precision in vulnerability localization.
@@
To address this, LOVA implements a learning approach that integrates sequential information by employing a Bi-LSTM model, which effectively captures these contextual dependencies: First, The identification of anomalies of VulnAttnMat is context-dependent. For example, when inputting three lines of code, we generate three corresponding VulnAttnMat: $A$, $B$, and $C$. If $A$ significantly deviates from $B$ and $C$, it suggests that $A$ may be the most suspicious. To effectively model this, a sequence-to-sequence approach is necessary, with LSTM (Sherstinsky, 2020) and Transformer being the most common models for such tasks. Second, given that the features of anomalies in our case are relatively distinct, a straightforward LSTM model is sufficient. However, we opted for a Bi-LSTM (Graves et al., 2013) to leverage its ability to capture bidirectional dependencies, as standard LSTM only processes information in one direction. This is important because the detection of anomalies is influenced by both the preceding and succeeding contexts, rather than solely relying on the prior context.
@@
where $N$ is the number of lines, $y_{i}$ is the true label of the $i$-th line, and $\hat{y}_{i}$ is the predicted probability of the line containing a vulnerability. During the evaluation phase, the flattened vectors of the VulnAttnMat are input into the model to compute suspicious scores for each line. These scores are subsequently ranked in descending order to inform the vulnerability localization process, with scores greater than 0.5 being considered vulnerable lines.
5. Evaluation
@@
    How does the design of LOVA impact the results (Ablation Study)?
5.1. Experiment Setup
Dataset: Our dataset, comprises vulnerability benchmarks from C/C++ projects (Big-Vul), Solidity smart contracts (SmartFix), and multi-language CVE records (CVEFixes). This allows for a comprehensive evaluation of LOVA’s performance across diverse contexts. Specifically:
  • Big-Vul (Fan et al., 2020): A C/C++ code vulnerability benchmark collected from 384 open-source GitHub projects. Big-Vul includes both vulnerable code snippets and their corresponding fixing commits. Vulnerabilities are annotated at function context level, categorizing the types of vulnerabilities and associated commits. Big-Vul contains a total of 11,823 vulnerable functions.
  • SmartFix (So and Oh, 2023): A Solidity code vulnerability benchmark collected from multiple sources. SmartFix covers five common types of smart contract vulnerabilities: Integer Overflow, Reentrancy, Ether Leak, Suicidal, and incorrect use of ‘tx.origin‘. Each vulnerability is annotated with its corresponding line numbers. SmartFix contains a total of 361 contract files.
  • CVEFixes (Bhandari et al., 2021): A multi-language code vulnerability benchmark that is automatically collected and curated from CVE records in the public U.S. National Vulnerability Database (NVD). CVEFixes collects both vulnerabilities and the commits that fix them, including 5,365 CVEs and 50,322 methods. CVEFixes encompasses multiple languages, such as C, C++, PHP, Python, Java, and JavaScript. From these, we selected the three most commonly used languages, C, Python, and Java, as benchmarks.
Big-Vul provides vulnerability data at the function context level, CVEFixes provides vulnerability data at both the function and file context levels, and SmartFix provides vulnerability data at the contract context level. To evaluate LOVA’s performance across different contexts, we conduct evaluations at function, file, and contract context levels. Additionally, we filter out contexts exceeding 4,000 tokens to comply with the input constraints of LLMs. In total, seven datasets were created, as shown in Table 1. Among them, the C datasets Big-Vul and CVEFixes-C each contain 1,000 instances. For Java, there are two datasets: CVEFixes-J-M and CVEFixes-J, corresponding to method-level and file-level contexts, respectively. In Python, the CVEFixes-P-M and CVEFixes-P datasets represent method-level and file-level contexts, respectively. Additionally, Solidity has one dataset, SmartFix, focusing on contract-level context.
Table 1. All datasets, C, J, and P, represent programs written in C, Java, and Python, respectively.